OpenStudy (anonymous):

a square based pyramid with side 18cm and height 12cm, calculate the surface area.

OpenStudy (anonymous):

a=2bs + b^2 a=2(18)(6√13) + (18)^2 a = 1102.799076 I searched the formula up online and i did that. idk what to do next.

OpenStudy (jdoe0001):

let us look at the bottom of the SQUARE BASE pyramid

OpenStudy (jdoe0001):

what do you think will be half-way to the center towards the "apothem"?

OpenStudy (anonymous):

9?

OpenStudy (jdoe0001):

right, now lemme stand up your pyramid, and let us find the so-called "slant height", using the pythagorean theorem

OpenStudy (jdoe0001):

keep in mind that \(\bf c^2 = a^2 + b^2\)

OpenStudy (anonymous):

c^2= a^2 + b^2 c^2 = 12^2 + 9^2 c^2 = 144 +81 c^2 = 225 √c^2= √225 c = 15

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

so now you have the slant height and the slant height, which is really the "nose of the triangular face" is also the height of that triangular face, and the area of a triangle is 1/2 (base)(height) you have 4 triangular faces, add them all 4 then add to that the Area of the squarish bottom and that's the area of the pyramid

OpenStudy (anonymous):

4(9) +18 = 54?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

area of one lateral face = 1/2 (9*15) =67.5 SA = 18 +4(67.5) SA =288?